OpenStudy (anonymous):

Simplify completely 25^3 -35x^2 +5 / -5x

OpenStudy (anonymous):



OpenStudy (jennychan12):

hello again. ahh, algebra... \[\frac{ 25x^3-35x+5 }{ -5x}\] split the terms \[\frac{ 25x^3 }{ -5x } +\frac{ 35x }{ 5 } -\frac{ 5 }{ 5x }\] then simpliy each term got it from here?

OpenStudy (anonymous):

Quotient : - 5x2 + 7x =x(-5x+7) Remainder : 5

OpenStudy (jennychan12):

\[-5x^2 +7x -\frac{ 1 }{ x }\] you could combine it into \[\frac{ -5x^3+7x^2-x }{ x }\]

OpenStudy (mayankdevnani):